Joe buys a set of books that cost $45.58. He gives the cashier three (3) $20 bills and three (3) pennies. How much change should Joe get back?
Answer by fractalier(6550)

3(20) + .03 - 45.58 =
60.03 - 45.58 =
$14.45 change